# Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model

### v2

@@ Abstract
Exact numerical results.—Previously, we carried out a comparison between simple analytical expressions depending on elementary functions, Eqs. (8) and (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model), and full analytical expressions, Eqs. (5) and (6), that still require the diagonalization of a single-particle random matrix $J_{ij}$. We now compare those analytical results with a numerical calculation of $C(t)$ by exact diagonalization combined with the quantum trajectory method [71, 69, 70] when the bath is turned on ($\mu\neq 0$). The fermionic quantum trajectory method is described in Appendix B. The need of a double average over quantum trajectories and disorder realizations for $\mu>0$ limits the sizes for which we can make a detailed comparison to $N\lesssim 20$.
@@
The results in Fig. 2 for $\mu=0$ show an excellent agreement between the compact analytical $C(t)$ Eq. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model), including $1/N+1/N^{2}$ contributions, and the exact diagonalization result. The small shift upward of the former is of order $1/N^{3}$ and therefore consistent with neglected terms in the expansion. Differences between the exact analytical $C(t)$ Eqs. (5), (6) and exact diagonalization results are barely noticeable. We stress that the comparison is parameter free and that the inclusion of $1/N^{2}$ corrections is essential for the observed level of agreement. For a finite $\mu$ (bottom plot), the numerical $C(t)$, using $200$ quantum trajectories, reproduces correctly the expected ${\rm e}^{-2\mu t}$ decay due to the bath, a feature that we derive analytically in Appendix B. Subleading features, like oscillations, are difficult to reproduce numerically because it would require a much larger number of disorder realizations. This illustrates the importance of obtaining analytic results to describe quantitatively the different stages of the quantum dynamics.
@@
Conclusions.—We studied information scrambling and the effect of Markovian dissipation in an integrable SYK model through an analytical calculation of the growth of quantum uncertainty characterized by OTOCs at all times scales. In the absence of dissipation, the asymptotic approach to the steady state is power-law with superimposed oscillations. At $t\sim\sqrt{N}$, the overall growth of scrambling stops and starts to decrease linearly in time. This unexpected change of trend has its origin in a subleading $1/N^{2}$ correction to the OTOC that becomes dominant in this region. We show analytically that this correction is nothing but the SFF of the random couplings $J_{ij}$. It becomes dominant for $t\gtrsim\sqrt{N}$ because the SFF is linear in time, whereas the rest of the time-dependent terms in the OTOC tend to zero in a power-law fashion. For $t\approx 2N$ (the Heisenberg time), the linear decrease of the uncertainty terminates, and the uncertainty reaches saturation though we still observe small oscillatory contributions whose amplitude decreases as a power law in time. The effect of the Markovian environment is an exponential decay of the growth of quantum uncertainty at a timescale inversely proportional to the coupling to the bath, so it will eventually dominate the approach to saturation. Oscillations in time are still occur but are a subleading effect in this limit. We believe that features like a power-law decay approach to saturation together with an oscillatory behavior of the OTOC are generic features in quantum many-body integrable systems. It would be interesting to explore the dynamics of the present SYK model employing more general jump operators so that the vectorized Liouvillian has random quartic terms in order to study whether the environment can induce quantum chaotic features in the dynamics.
@@
Acknowledgments.—AMGG thanks Klaus Mølmer for illuminating correspondence. AMGG, JPZ, and CL acknowledge support from the National Natural Science Foundation of China (NSFC): Individual Grant No. 12374138, Research Fund for International Senior Scientists No. 12350710180, and National Key R&\&&D Program of China (Project ID: 2019YFA0308603). AMGG acknowledges support from a Shanghai talent program. LS was supported by a Research Fellowship from the Royal Commission for the Exhibition of 1851. JJMV is supported in part by US DOE Grant No. DE-FAG-88FR40388.
@@ Appendix A Finite-temperature results
Finite-temperature steady state.—By an orthogonal transformation of the Majorana fermions, $\widetilde{\chi}_{i}=S_{ij}\chi_{j}$, the $q=2$ Hamiltonian of the SYK model in Eq. (1) can be rewritten as [72]
@@
Factorization of the $\mu$-dependence.—The time evolution of $c_{p}$ and $c_{p}^{\dagger}$ is given by the adjoint Lindblad operator [59],
@@
Because the original Majoranas are linearly related to the $\widetilde{\chi}_{k}$ this equation also holds for the $\chi_{k}$. This shows that the $\mu$-dependence also factorizes at finite temperature.
Figure 3: $C(t)$ with $t$ in units of $J=1$ for $\beta=5$. Top: $\mu=0$. The black curve stands for the semiclassical expression Eq. (A) with $C_{N}$ given by the time independent part of the coefficient of $\exp[-\mu t]$ in Eq. (A). The blue ($N=20$), red ($N=200$), and green ($N=2000$) curves stand for the exact analytic $C(t)$, Eqs. (5) and (6) with $C_{N}=0.06,0.044,0.043$ for $N=20,200,2000$, respectively. The initial approach to the steady state is still power-law, so thermal effects do not lead to qualitative changes. Bottom: $\mu=0.15,\beta=5$. For $t\gtrsim 1/2\mu$ the decay is exponential and fully controlled by the Markovian bath.
Dynamics of $C(t)$.—At finite $\beta$, the leading $1/N$ expression for $C(t)$ is [recall that at $\beta=0$ we have Eq. (7)]:
@@
[…]
@@
The one-dimensional integrals in Eq. (A) cannot be evaluated exactly but an asymptotic expression valid in the limit $t\gg\beta$ with $\beta$ and $\mu$ fixed is available. The most salient finite-$\beta$ effect in $C(t)$ is to induce, for intermediate times, an exponential decay $\sim e^{-\beta\pi t}$. For longer times, the approach to the steady state is power-law $\sim 1/t^{3}$ as in the $\beta=0$ case. At finite $\mu$, the asymptotic decay is exponential and $\beta$-independent, with oscillations still superimposed. Therefore, thermal effects are mostly quantitative not qualitative. This is confirmed by an explicit comparison, depicted in Fig. 3, between the $1/N$ prediction Eq. (A) and the exact $C(t)$ Eqs. (5) and (6) for $\beta=5$.
Regarding $1/N^{2}$ corrections, an analogous expression to Eq. (Anatomy of information scrambling and decoherence in the integrable Sachdev-Ye-Kitaev model) could also be worked out at finite $\beta$, although it would be rather cumbersome. Moreover, the results of Fig. 3 show that thermal effects do not induce quantitative changes in $C(t)$.
@@ Appendix B Fermionic Quantum Trajectory Method
In this appendix, we first review the method of quantum trajectories [71, 69, 70], originally proposed in the context of bosonic systems and extend it to the fermionic systems discussed in this paper. We consider an open quantum system where the system degrees of freedom consist of $N$ bosonic fields $a_{i},i=1,\ldots,N$ with $[a_{i},a_{j}^{\dagger}]=\delta_{ij}$. When coupled to a Markovian bath, the reduced density matrix $\rho$ for the system evolves under the following master equation
@@
$\frac{dG_{ij}}{dt}={\rm i}\langle[H,a_{i}](t)\,a_{j}\rangle-\mu NG_{ij}(t)~+\mu\sum_{n}\langle(L^{\dagger}_{n}a_{i}L_{n})(t)\,a_{j}\rangle.$ (21)
@@
In order to extend these results to fermonic systems such as the ones considered in the main text, a central issue is the anti-commutation relations of the fermionic field operators that replace the commutation relations assumed in the analysis above. When the jump operators and the fermion operators anti-commute, the quantum master equation (20) governing the evolution of the reduced density matrix can no longer be directly applied to the $n$-point function. Instead, as is shown in Appendix B of Ref. [63], the adjoint Lindblad operator acts on the Green’s function with an extra minus sign. More specifically, using the notations of the main text where $\chi_{i},i=1,\ldots,N$ are $N$ Majorana fermions with anti-commutation relations $\{\chi_{i},\chi_{j}\}=\delta_{ij}$, Eq. (21) should be replaced with
$\frac{dG_{ij}}{dt}={\rm i}\langle[H,\chi_{i}](t)\,\chi_{j}\rangle-\mu NG_{ij}(t)~-\mu\sum_{n}\langle(L^{\dagger}_{n}\chi_{i}L_{n})(t)\,\chi_{j}\rangle.$ (25)
@@
$\widetilde{L}^{\dagger}_{n}\chi_{i}\widetilde{L}_{n}=L^{\dagger}_{n}\gamma_{c}\chi_{i}\gamma_{c}L_{n}=-L^{\dagger}_{n}\chi_{i}L_{n}.$ (26)
@@
The $\mu$-dependence of this time evolution can be worked out analytically when we consider a single time step. Since at every time step we have a probability $\mu N\delta t/2$ for a jump operator and a probability $(1-\mu N\delta t/2)$ for a Hamiltonian evolution, we have
$\frac{1}{M}\sum_{q=1}^{M}\widetilde{U}^{\dagger}_{q}(\delta t)\chi_{i}\widetilde{U}_{q}(\delta t)=\left(1-\frac{\mu N\delta t}{2}\right)~e^{iH\delta t}\chi_{i}~e^{-iH\delta t}$ (28)
$+$ $\frac{\mu N\delta t}{2}\frac{(N-2)}{N}\chi_{i}=e^{iH\delta t}\chi_{i}~e^{-iH\delta t}-\mu\delta t\chi_{i},$
where we have used $\widetilde{L}_{n}\chi_{i}\widetilde{L}_{n}=\chi_{i}$ for $i\neq n$ and $\widetilde{L}_{i}\chi_{i}\widetilde{L}_{i}=-\chi_{i}$. For $q=2$ and to linear order in $\delta t$, this implies that $\chi_{i}$ satisfies the adjoint master equation Eq. (18), implying that for $\mu t\ll N$ the $\mu$-dependence factorizes as
@@
We emphasize that the additional minus sign in the master equation was essential to obtain the correct $\mu$-dependence [53]. Had we not included it, we would instead obtain a factor $e^{-\mu Nt}$, which is what one expects for a bosonic system.
The procedure for the four-point function considered in the main text follows by replacing both time evolutions by the right-hand side of Eq. (28) using independent trajectories. From the argument above, it is clear that the $\mu$-dependence factorizes for the four-point function as well.

### v3

@@ IV Exact numerical results
Previously, we carried out a comparison between simple analytical expressions depending on elementary functions, Eqs. (15) and (III), and full analytical expressions, Eqs. (12) and (13), that still require the diagonalization of a single-particle random matrix $J_{ij}$. We now compare those analytical results with a numerical calculation of $C(t)$ by exact diagonalization combined with the quantum trajectory method [77, 75, 76] when the bath is turned on ($\mu\neq 0$). The fermionic quantum trajectory method is described in Appendix B. The need of a double average over quantum trajectories and disorder realizations for $\mu>0$ limits the sizes for which we can make a detailed comparison to $N\lesssim 20$.
@@
The results in Fig. 2 for $\mu=0$ show an excellent agreement between the compact analytical $C(t)$ Eq. (III), including $1/N+1/N^{2}$ contributions, and the exact diagonalization result. The small shift upward of the former is of order $1/N^{3}$ and therefore consistent with neglected terms in the expansion. Differences between the exact analytical $C(t)$ Eqs. (12), (13) and exact diagonalization results are barely noticeable. We stress that the comparison is parameter free and that the inclusion of $1/N^{2}$ corrections is essential for the observed level of agreement. For a finite $\mu$ (bottom plot), the numerical $C(t)$, using $200$ quantum trajectories, reproduces correctly the expected ${\rm e}^{-2\mu t}$ decay due to the bath, a feature that we derive analytically in Appendix B. Subleading features, like oscillations, are difficult to reproduce numerically because it would require a much larger number of disorder realizations. This illustrates the importance of obtaining analytic results to describe quantitatively the different stages of the quantum dynamics.
V Conclusions
@@
We studied information scrambling and the effect of Markovian dissipation in an integrable SYK model through an analytical calculation of the growth of quantum uncertainty characterized by OTOCs at all times scales. In the absence of dissipation, the asymptotic approach to the steady state is power-law with superimposed oscillations. At $t\sim\sqrt{N}$, the overall growth of scrambling stops and starts to decrease linearly in time. This unexpected change of trend has its origin in a subleading $1/N^{2}$ correction to the OTOC that becomes dominant in this region. We show analytically that this correction is nothing but the SFF of the random couplings $J_{ij}$. It becomes dominant for $t\gtrsim\sqrt{N}$ because the SFF is linear in time, whereas the rest of the time-dependent terms in the OTOC tend to zero in a power-law fashion. For $t\approx 2N$ (the Heisenberg time), the linear decrease of the uncertainty terminates, and the uncertainty reaches saturation though we still observe small oscillatory contributions whose amplitude decreases as a power law in time. The effect of the Markovian environment is an exponential decay of the growth of quantum uncertainty at a timescale inversely proportional to the coupling to the bath, so it will eventually dominate the approach to saturation. We believe that features like a power-law decay approach to saturation and a local maximum of the quantum uncertainty due to the discreteness of the spectrum are generic features in quantum many-body integrable systems. It would be interesting to explore the dynamics of the present SYK model employing more general jump operators so that the vectorized Liouvillian has random quartic terms in order to study whether the environment can induce quantum chaotic features in the dynamics.
Acknowledgements.
@@
AMGG thanks Klaus Mølmer for illuminating correspondence. AMGG, JPZ, and CL acknowledge support from the National Natural Science Foundation of China (NSFC): Individual Grant No. 12374138, Research Fund for International Senior Scientists No. 12350710180, and National Key R&\&D Program of China (Project ID: 2019YFA0308603). AMGG acknowledges support from a Shanghai talent program. LS was supported by a Research Fellowship from the Royal Commission for the Exhibition of 1851. JJMV is supported in part by US DOE Grant No. DE-FAG-88FR40388.
@@
A.1 Finite-temperature steady state
By an orthogonal transformation of the Majorana fermions, $\widetilde{\chi}_{i}=S_{ij}\chi_{j}$, the $q=2$ Hamiltonian of the SYK model in Eq. (1) can be rewritten as [78]
@@ A.1 Finite-temperature steady state
A.2 Factorization of the $\mu$ dependence
The time evolution of $c_{p}$ and $c_{p}^{\dagger}$ is given by the adjoint Lindblad operator [59],
@@ A.2 Factorization of the $\mu$ dependence
Because the original Majoranas are linearly related to the $\widetilde{\chi}_{k}$ this equation also holds for the $\chi_{k}$. This shows that the $\mu$ dependence also factorizes at finite temperature.
Figure 3: $C(t)$ with $t$ in units of $J=1$ for $\beta=5$. Top: $\mu=0$. The black curve stands for the semiclassical expression Eq. (A.3) with $C_{N}$ given by the time independent part of the coefficient of $\exp[-\mu t]$ in Eq. (A.3). The blue ($N=20$), red ($N=200$), and green ($N=2000$) curves stand for the exact analytic $C(t)$, Eqs. (12) and (13) with $C_{N}=0.06,0.044,0.043$ for $N=20,200,2000$, respectively. The initial approach to the steady state is still power-law, so thermal effects do not lead to qualitative changes. Bottom: $\mu=0.15,\beta=5$. For $t\gtrsim 1/2\mu$ the decay is exponential and fully controlled by the Markovian bath.
A.3 Dynamics of $C(t)$
At finite $\beta$, the leading $1/N$ expression for $C(t)$ is [recall that at $\beta=0$ we have Eq. (14)]:
@@ A.3 Dynamics of $C(t)$
$-3\left(\int_{-2}^{2}d\lambda\rho(\lambda)\frac{\cos(\lambda t)\cosh\frac{\beta\lambda}{4}}{\cosh\frac{\beta\lambda}{2}}\right)^{2}$
@@
The one-dimensional integrals in Eq. (A.3) cannot be evaluated exactly but an asymptotic expression valid in the limit $t\gg\beta$ with $\beta$ and $\mu$ fixed is available. The most salient finite-$\beta$ effect in $C(t)$ is to induce, for intermediate times, an exponential decay $\sim e^{-\beta\pi t}$. For longer times, the approach to the steady state is power-law $\sim 1/t^{3}$ as in the $\beta=0$ case. At finite $\mu$, the asymptotic decay is exponential and $\beta$-independent, with oscillations still superimposed. Therefore, thermal effects are mostly quantitative not qualitative. This is confirmed by an explicit comparison, depicted in Fig. 3, between the $1/N$ prediction Eq. (A.3) and the exact $C(t)$ Eqs. (12) and (13) for $\beta=5$.
Regarding $1/N^{2}$ corrections, an analogous expression to Eq. (III) could also be worked out at finite $\beta$, although it would be rather cumbersome. Moreover, the results of Fig. 3 show that thermal effects do not induce quantitative changes in $C(t)$.
@@
In this appendix, we first review the method of quantum trajectories [77, 75, 76], originally proposed in the context of bosonic systems and extend it to the fermionic systems discussed in this paper. We consider an open quantum system where the system degrees of freedom consist of $N$ bosonic fields $a_{i},i=1,\ldots,N$ with $a_{i},a_{j}^{\dagger}]=\delta_{ij}$. When coupled to a Markovian bath, the reduced density matrix $\rho$ for the system evolves under the following master equation
@@
$\frac{dG_{ij}}{dt}={\rm i}\langle[H,a_{i}](t)\,a_{j}\rangle-\mu NG_{ij}(t) +\mu\sum_{n}\langle(L^{\dagger}_{n}a_{i}L_{n})(t)\,a_{j}\rangle.$ (29)
@@
In order to extend these results to fermonic systems such as the ones considered in the main text, a central issue is the anti-commutation relations of the fermionic field operators that replace the commutation relations assumed in the analysis above. When the jump operators and the fermion operators anti-commute, the quantum master equation (28) governing the evolution of the reduced density matrix can no longer be directly applied to the $n$-point function. Instead, as is shown in Appendix B of Ref. [63], the adjoint Lindblad operator acts on the Green’s function with an extra minus sign. More specifically, using the notations of the main text where $\chi_{i},i=1,\ldots,N$ are $N$ Majorana fermions with anti-commutation relations $\{\chi_{i},\chi_{j}\}=\delta_{ij}$, Eq. (29) should be replaced with ($G_{ij}(t)=\langle\chi_{i}(t)\chi_{j}\rangle$)
$\frac{dG_{ij}}{dt}={\rm i}\langle[H,\chi_{i}](t)\,\chi_{j}\rangle-\mu NG_{ij}(t) -\mu\sum_{n}\langle(L^{\dagger}_{n}\chi_{i}L_{n})(t)\,\chi_{j}\rangle.$ (33)
@@
$\widetilde{L}^{\dagger}_{n}\chi_{i}(t)\widetilde{L}_{n}=L^{\dagger}_{n}\gamma_{c}\chi_{i}(t)\gamma_{c}L_{n}=-L^{\dagger}_{n}\chi_{i}(t)L_{n}.$ (34)
@@
The $\mu$ dependence of this time evolution can be worked out analytically when we consider a single time step. Since at every time step we have a probability $\mu N\delta t/2$ for a jump operator and a probability $(1-\mu N\delta t/2)$ for a Hamiltonian evolution, we have
$\hskip-14.22636pt\frac{1}{M}\sum_{q=1}^{M}\widetilde{U}^{\dagger}_{q}(\delta t)\chi_{i}(t)\widetilde{U}_{q}(\delta t)=\left(1-\frac{\mu N\delta t}{2}\right) e^{iH\delta t}\chi_{i}(t) e^{-iH\delta t}$ (36)
$+$ $\frac{\mu N\delta t}{2}\frac{(N-2)}{N}\chi_{i}(t)=e^{iH\delta t}\chi_{i}(t) e^{-iH\delta t}-\mu\delta t\chi_{i}(t),$
where we have used $\widetilde{L}_{n}\chi_{i}\widetilde{L}_{n}=\chi_{i}$ for $i\neq n$ and $\widetilde{L}_{i}\chi_{i}\widetilde{L}_{i}=-\chi_{i}$. For $q=2$ and to linear order in $\delta t$, this implies that $\chi_{i}$ satisfies the adjoint master equation Eq. (26), implying that for $\mu t\ll N$ the $\mu$ dependence factorizes as
@@
We emphasize that the additional minus sign in the master equation was essential to obtain the correct $\mu$ dependence [53]. Had we not included it, we would instead obtain a factor $e^{-\mu Nt}$, which is what one expects for a bosonic system.
The procedure for the four-point function considered in the main text follows by replacing both time evolutions by the right-hand side of Eq. (36) using independent trajectories. From the argument above, it is clear that the $\mu$ dependence factorizes for the four-point function as well.